write a quadratic equation that has the solutions -10i and 10i
thank you

Answer by jai_kos(139)

Given the solutions are -10i and +10i
Let it be the solutions for x
x = -10i or x = 10i
(x +10i)= 0 or (x -10i) =0
multiply both the equations we get
(x+10i)(x-10i) = 0
Using the foil method, we expand the above.
x^2 +10ix -10ix +10i* -10i = 0
x^2 -100(i)^2 = 0
Because (i)^2 = -1, we substitute above, we get
x^2 -100(-1) = 0
x^2 +100 = 0
Therefore the quadratic equation which has the root -10i and 10i ia x^2 + 100 = 0